OpenStudy (anonymous):

Jack's father let Jack use a section of the backyard that had dimensions of 10 ft and 12 ft for his garden. He wanted to put a fence around the outside of the garden. How long should the piece of fencing he needs to buy be?

OpenStudy (anonymous):

calculate perimeter p=2(l+b)...use this formula

Directrix (directrix):

@OneKoi Assuming the garden is rectangular, the perimeter of the garden is the perimeter of the fencing. P = 2*10 + 2*12 = 20 +24 = 44 feet. Note: Opposite sides of a rectangle have equal lengths